# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v1

@@ 5.1 Experimental Setup
We trained Multi-Layer Perceptions (MLP) Neural Networks from scikit-learn (https://scikit-learn.org/) and Gradient-boosted Tree Ensemble classifiers (Random Forests-based) from the XGBoost gradient boosting library [21] as the black-box models used throughout the experiments. Logistic Regressions and SVMs could also be viable alternatives, but we focused on Neural Networks and Random Forests due to their wide adoption in Machine Learning. The models were trained with a dataset split of $80\%$ for training and $20\%$ for testing. Such division was achieved using the $train\_test\_split$ method from scikit-learn, with a consistent shuffling seed across all datasets. We conducted a Grid Search to determine the hyperparameters that optimize the models’ performance.
For the MLPs, we initialized the hyperparameters search using powers of $2$ for the number of neurons in the hidden layers, following standard practices in this context [60]. We employed the ReLU activation function, Stochastic Gradient Descent as the optimizer and the log-loss function. We evaluated alternative activation functions and optimizers, but they did not yield gains. The Tree Ensemble model’s hyperparameters were defined by specifying ranges for the number of decision tree estimators and the maximum depth of each estimator, utilizing cross-entropy loss as the evaluation metric for classification.
We selected three different MLP models from the hyperparameters tunning: a three and a five hidden layers neural network holding $64$ neuron units per layer (3H-NN and 5H-64-NN, respectively) and a five-layer neural network with $[64,128,128,128,64]$ neurons in each layer (5H-128-NN). All the neural network models use a learning rate of $0.01$, alpha $0.0001$, and maximum training epochs of $500$. The selected neural networks achieved better performances considering our classification tasks. We based our architectures on previous works [9, 15], with extra neural units and additional layers significantly increasing the training time without noticeably increasing performance.
Following our hyperparameters tunning, the Tree Ensemble classifiers (XRFC) have $500$ estimators, a maximum depth of $6$, a learning rate of $0.01$, and a gamma equal to $1$. All the other hyperparameters are the libraries’ default for Neural Networks and Tree Ensembles. Those models are used as the base black boxes for comparing T-Explainer against SHAP [40], LIME [48], and three Gradient-based methods [32] such as Integrated Gradients, Input $\times$ Gradient, and DeepLIFT. Specifically, we used the SHAP explainer for neural networks and the TreeSHAP explainer [41] for tree-based classifiers. Although SHAP, LIME, and Gradient-based methods were proposed a few years back, we benchmark T-Explainer with them because they continue to be the most widely used feature attribution explainers in research and practice [4].
@@ 5.2 Synthetic Data
We generate two different synthetic datasets comprising $1000$ instances each. The first is a 4-dimensional (4-FT) dataset where each instance $\mathbf{x}$ is generated as follows. We distribute the target label $\mathbf{y}\in[0,1]$ equally across each dataset half; thus, each class has $500$ instances. Conditioned to the value of $\mathbf{y}$, we sample the instance $\mathbf{x}$ as $x_{1:2}\sim\mathcal{N}(\mu_{y},\,\Sigma_{y})$. We choose $\mu_{0}=[0,0]^{\mathrm{T}}$ and $\mu_{1}=[-2,2]^{\mathrm{T}}$, $\Sigma_{0}=[[1,1],[-1,1]]^{\mathrm{T}}$ and $\Sigma_{1}=\mathbf{I}$, where $\mu_{0}$, $\Sigma_{0}$ and $\mu_{1}$, $\Sigma_{1}$ denote the means and covariance matrices of the Gaussian distributions associated with instances in classes 00 and $1$ respectively. The features $x_{1:2}$ are called $core\_1$ and $core\_2$. Random values are assigned to the features $x_{3:4}$ ($noise\_1$ and $noise\_2$). Such configuration results in a dataset holding two predictive (important) and two random noise (non-important) features with a small mixture area between the important features to introduce a degree of complexity to the predictive process.
@@
The second synthetic dataset is more robust. It is a 20-dimensional (20-FT) dataset created using the OpenXAI synthetic data generation tool, whose algorithm is described in Agarwal et al. [4]. According to the authors, the algorithm ensures the creation of a dataset that encapsulates feature dependencies and clear local neighborhoods, key properties to guarantee the explanations derived from this synthetic dataset remain consistent with the behavior of the models trained on such data.
@@
We used the XRFC and 3H-NN models over the synthetic datasets experiments. The XRFC classifier achieved $96\%$ accuracy on the 4-FT data and $83.5\%$ accuracy on the 20-FT synthetic dataset, while the 3H-NN model achieved $97.5\%$ accuracy on the 4-FT data and $83.5\%$ accuracy on the 20-FT dataset. Note that we are not basing our benchmark experiments on synthetic datasets with massive amounts of instances. We know it is necessary to use a reasonable amount of data to train and test machine learning models. According to Aas et al. [1], well-known feature importance methods become unstable in tasks with more than ten dimensions. In this sense, data dimensionality is more critical to assessing the stability of feature importance methods than the number of instances. Thus, we generated synthetic datasets with enough instances ($1000$ for each dataset) to train our models and assess the XAI methods.
Figures 2(a) and 2(b) present explanations generated by TreeSHAP and T-Explainer for the same instance predicted with the XRFC model. As expected, TreeSHAP attributed the most significant importance to feature $core\_1$ (followed by $core\_2$). However, TreeSHAP attributed a negative signal to $core\_1$, pushing the additive reconstruction (see Equation 5) of TreeSHAP to $-0.013$, while the true prediction for this instance is $0.0134$. Thus, the TreeSHAP explanation indicates that the instance belongs to the opposite class, disagreeing with the model prediction (and also the instance label) and violating the Local Accuracy property. T-Explainer also points to $core\_1$ (and $core\_2$) as the most important feature but with a correct positive signal.
Table 1 provides a broad view of the scenario illustrated in Figure 2. We used the FAE metric (considering $0.01$ as the tolerance threshold) to assess Local Accuracy preservation by T-Explainer, SHAP, and TreeSHAP when applied to the XRFC model on the 4-FT dataset. TreeSHAP, the model-specific version of SHAP, presented a faithful additivity rate of $0.507$ of its explanations. T-Explainer preserved the additivity property with a $0.425$ rate, slightly lower than TreeSHAP. However, the T-Explainer was far better than the SHAP model-agnostic version of the SHAP explainer framework. In contrast, Table 2 shows that the T-Explainer outperforms all the other additive methods regarding Local Accuracy preservation on the 20-FT dataset holding a $0.862$ preservation rate.
(a) TreeSHAP feature importances.
(b) T-Explainer feature importances.
Figure 2: TreeSHAP and T-Explainer importances of the same instance classified with XRFC on the 4-FT dataset.
XRFC FAE
XAI Mean Ratio
T-Explainer 0.0178 0.425
SHAP 0.5072 0.007
TreeSHAP 0.0186 0.507
Table 1: T-Explainer, SHAP, and TreeSHAP Local Accuracy when explaining the XRFC predictions on the 4-FT dataset.
XRFC FAE
XAI Mean Ratio
T-Explainer 0.0053 0.862
SHAP 0.4601 0.319
TreeSHAP 0.0103 0.620
Table 2: T-Explainer, SHAP, and TreeSHAP Local Accuracy when explaining the XRFC predictions on the 20-FT dataset.
@@
Table 3: Stability of XAI methods explaining the XRFC predictions on the 4-FT synthetic dataset. T-Exp is the T-Explainer, and Max refers to the maximum value.
@@
Table 4: Stability of XAI methods explaining the 3H-NN predictions on the 4-FT dataset. I-Grad and I×$\times$×Grad are Integrated Gradients and Input $\times$ Gradient, respectively.
Tables 5 and 6 depict the results from the XRFC and 3H-NN models on the 20-FT dataset. From Table 5, we can see that T-Explainer is superior to TreeSHAP and LIME in the RIS and ROS metrics for both maximum and mean values. TreeSHAP reached a close performance regarding maximum ROS, but T-Explainer is considerably better on the average ROS. Similar results can be observed in Table 6, where T-Explainer beat all the other explainers in terms of RIS (3×3$\times$3 × better than Integrated Gradients), with slightly better performance than Integrated Gradients in the mean ROS but outperforming all the other explainers in maximum ROS. SHAP was the most unstable method, which can be justified in the context of the 20-FT dataset because, in high-dimensional spaces, SHAP uses a random sampling algorithm rather than the deterministic one, leading to unstable explanations [6]. Moreover, SHAP is also prone to suffer from extrapolations [30, 31].
@@
In general, Tables 3, 4, 5, and 6 show that T-Explainer is robust regarding different stability metrics, being among the best performance methods, especially when data dimensionality is high.
The T-Explainer’s performance when dealing with the XRFC model, which is a non-differentiable model, is also remarkable. Therefore, although not theoretically supported, the T-Explainer might perform well even when applied to generate explanations for non-differentiable machine learning models.
5.3 Real Data
This Section extends the evaluations and comparisons by applying real data from different domains. Specifically, we run experiments using four well-known real datasets with distinct properties regarding dimensionality, the presence of categorical attributes, and size.
@@ 5.3 Real Data
The Home Equity Line of Credit (HELOC) dataset provided by FICO [26] consists of financial attributes from anonymized applications for home equity lines of credit submitted by $9,871$ real homeowners. The task in the HELOC dataset is to predict whether an applicant has a good or bad risk of repaying their HELOC account within two years. The largest dataset in our study is HIGGS [9], which contains $28$ features about simulated collision events to distinguish between Higgs bosons signals and a background process. The original HIGGS dataset contains $11$ million instances, but we used the $98,050$ instances version available at OpenML [64]. Table 7 summarizes the real datasets.
@@
The experiments with real data have been conducted using only the neural network models, as some of the explanation models with which T-Explainer is compared are specifically designed for neural networks. Moreover, the neural network models are differentiable, thus meeting the theoretical requirements that support the T-Explainer technique. We further discuss tree-based models in Section 7.
For the experiments with Banknote Authentication and German Credit datasets, we selected the 3H-NN model. From Table 8, one can see that T-Explainer is less stable than Integrated Gradients, but it is the second-best method on average for RIS and ROS (closely tied with Integrated Gradients for RES). Table 9 presents the stability tests using the German Credit data. The original version of the dataset has $20$ numerical and categorical attributes. Still, it is practically impossible to directly apply it in a machine learning task due to its complex categorization system.
We pre-processed the German Credit data to clean it, resulting in a reduced version holding nine features ($4$ numerical and $5$ categorical features, see Table 7). As shown in Table 9, T-Explainer performed as the most stable explainer regarding RIS and maximum ROS perturbations, being the second best for the mean ROS.
@@
Tables 8 and 9 indicate that gradient-based methods are more stable to input/output perturbations than SHAP and, in some cases, also than LIME. As one can notice, T-Explainer competes quite well with other gradient-based methods.
HELOC and HIGGS are more robust datasets in terms of dimensionality and size. We selected the 3H-NN and the five-layer neural network models to compare T-Explainer with the other explainers on those datasets. Specifically, we trained the 3H-NN and 5H-128-NN classifiers on HELOC, achieving $71.34\%$ and $73.27\%$ accuracy, respectively. Despite the accuracy values being lower than the ones we achieved before, those performances are in line with the results reported in the literature [15].
Table 10 shows that Integrated Gradients is the most stable explainer for almost all metrics (behind LIME only for mean RIS but with very close numbers). However, we highlight the performance of T-Explainer, especially regarding the metrics’ means. Only Integrated Gradients, LIME, and T-Explainer kept the mean RIS under ten units, which can be taken as virtually the same mean stability to input perturbations. Similar results are observed in mean ROS, where Integrated Gradients, LIME, and T-Explainer achieved the smallest values, being T-Explainer the second best. Table 11 shows the explainers’ stability when applied to the five-layer neural network. In such case, T-Explainer is the most stable method, with Integrated Gradients being the second best for mean RIS and LIME as the second best for mean ROS. Only T-Explainer and Integrated Gradients matched RES with the highest levels of stability.
@@
We trained the 3H-NN and 5H-64-NN classifiers on the HIGGS data, which achieved $69.58\%$ and $66.83\%$ accuracy, respectively, close to the values reported in previous works [15]. We included a set of benchmarks using a five-layer neural network here to keep some similarity with the architectures proposed in Baldi et al. [9], which explored the use of deep networks on HIGGS. Tables 12 and 13 show the good stability performances of gradient-based feature importance explainers, with T-Explainer outperforming them regarding RIS perturbations while being quite competitive regarding ROS and RES metrics. Those results demonstrate T-Explainer’s ability to generate consistent explanations through different levels of model complexity and data dimensionality. Note that SHAP is more unstable than LIME, with SHAP’s claimed advantageous performance predominantly seen in low-dimensional datasets [6].
@@
Finally, the experiments above show that T-Explainer generally performs well, clearly outperforming well-known model-agnostic methods such as SHAP (and LIME for most tests). Our method presented competitive results in setups differing from Neural Networks, proving its agnostic capability. Moreover, T-Explainer also turned out to be quite competitive with model-specific techniques such as the gradient-based ones, thus being a new and valuable alternative method to explain black-box models’ predictions.

### v2

@@ 5.1 Experimental Setup
We trained Neural Networks from scikit-learn (https://scikit-learn.org/) and Gradient-boosted Tree Ensemble classifiers (Random Forests-based) from the XGBoost gradient boosting library [20] as the black-box models used throughout the experiments. Logistic Regressions and SVMs could also be viable alternatives, but we focused on Neural Networks and Random Forests due to their wide adoption in Machine Learning. The models were trained with a dataset split of $80\%$ for training and $20\%$ for testing. Such division was achieved using the $train\_test\_split$ method from scikit-learn, with a consistent shuffling seed across all datasets. We conducted a Grid Search to determine the hyperparameters that optimize the models’ performance.
For the Neural Networks, we initialized the hyperparameters search using powers of $2$ for the number of neurons in the hidden layers, following standard practices in this context [61]. We employed the ReLU activation function, Stochastic Gradient Descent as the optimizer and the log-loss function. We evaluated alternative activation functions and optimizers, but they did not yield gains. The Tree Ensemble model’s hyperparameters were defined by specifying ranges for the number of decision tree estimators and the maximum depth of each estimator, utilizing cross-entropy loss as the evaluation metric for classification.
We selected three different Neural Network models from the hyperparameter tuning: a three- and a five-hidden-layer neural network holding $64$ neuron units per layer (3H-NN and 5H-64-NN, respectively) and a five-layer neural network with $[64,128,128,128,64]$ neurons in each layer (5H-128-NN). All the neural network models use a learning rate of $0.01$, alpha $0.0001$, and maximum training epochs of $500$. The selected neural networks achieved better performances considering our classification tasks. We based our architectures on previous works [10, 16], with extra neural units and additional layers significantly increasing the training time without noticeably increasing performance.
Following our hyperparameters tunning, the Tree Ensemble classifiers (XRFC) have $500$ estimators, a maximum depth of $6$, a learning rate of $0.01$, and a gamma equal to $1$. All the other hyperparameters are the libraries’ default for Neural Networks and Tree Ensembles. Those models are used as the base black boxes for comparing T-Explainer against SHAP [39], LIME [47], and three gradient-based methods from the Captum library [31] – Integrated Gradients, Input $\times$ Gradient, and DeepLIFT. Specifically, we used the SHAP explainer for neural networks and the TreeSHAP explainer [40] for tree-based classifiers. Although SHAP, LIME, and Gradient-based methods were proposed a few years back, we benchmark T-Explainer with them because they continue to be the most widely used feature attribution explainers in research and practice [4].
@@ 5.2 Synthetic Data
We generated two different synthetic datasets comprising $1,000$ instances each. The first is a 4-dimensional (4-FT) dataset where each instance $\mathbf{x}$ is generated as follows. We distribute the target label $\mathbf{y}\in[0,1]$ equally across each dataset half; thus, each class has $500$ instances. Conditioned to the value of $\mathbf{y}$, we sample the instance $\mathbf{x}$ as $x_{1:2}\sim\mathcal{N}(\mu_{y},\,\Sigma_{y})$. We choose $\mu_{0}=[0,0]^{\mathrm{T}}$ and $\mu_{1}=[-2,2]^{\mathrm{T}}$, $\Sigma_{0}=[[1,1],[-1,1]]^{\mathrm{T}}$ and $\Sigma_{1}=\mathbf{I}$, where $\mu_{0}$, $\Sigma_{0}$ and $\mu_{1}$, $\Sigma_{1}$ denote the means and covariance matrices of the Gaussian distributions associated with instances in classes 00 and $1$ respectively. The features $x_{1:2}$ are called $core\_1$ and $core\_2$. Random values are assigned to the features $x_{3:4}$ ($noise\_1$ and $noise\_2$). Such configuration results in a dataset holding two predictive (important) and two random noise (non-important) features with a small mixture area between the important features to introduce some degree of complexity.
@@
The second synthetic dataset is more robust. It is a 20-dimensional (20-FT) dataset created using the OpenXAI synthetic data generation tool, whose algorithm is described in Agarwal et al. [4]. According to the authors, the algorithm ensures the creation of a dataset that encapsulates feature dependencies and clear local neighborhoods, key properties to guarantee the explanations derived from this synthetic dataset remain consistent with the behavior of the models trained on such data. Using controlled synthetic data for XAI checking meets the correctness dimension of explainability evaluation [42].
We trained the XRFC and 3H-NN models for the synthetic datasets experiments. The XRFC classifier achieved $96\%$ accuracy on the 4-FT data and $83.5\%$ accuracy on the 20-FT data, while the 3H-NN model achieved $97.5\%$ accuracy on the 4-FT data and $83.5\%$ accuracy on the 20-FT synthetic dataset. Evaluating explanation methods on accurate models is good practice in controlled settings based on synthetic data [42]. The higher the model’s accuracy level, the more it can be assumed that the model adhered to the reasoning designed in the data.
@@
Note that we are not basing our benchmark experiments on synthetic datasets with massive amounts of instances. We know it is necessary to use a reasonable amount of data to train and test machine learning models. According to Aas et al. [1], well-known feature importance methods become unstable in tasks with more than ten dimensions. In this sense, data dimensionality is more critical to assessing the stability of feature importance methods than the number of instances. Thus, we generated synthetic datasets with enough instances ($1,000$ for each dataset) to train our models and assess the XAI methods.
@@
Table 1: Stability of XAI methods explaining the XRFC predictions on the 4-FT synthetic dataset. T-Exp denotes T-Explainer and Max refers to maximum values.
@@
Table 2: Stability of XAI methods explaining the 3H-NN predictions on the 4-FT dataset. I-Grad and I×$\times$×Grad denote Integrated Gradients and Input $\times$ Gradient, respectively.
Table 3 presents the PGI results of the explainers applied to the 3H-NN predictions on the 4-FT data. T-Explainer explanations are considerably more faithful than the other methods in identifying the most important feature for predictions (Top 1 column), i.e., T-Explainer’s explanations demonstrated the closest behavior to what the underlying model learned as the most important feature. When the task considered more features, the second and third most important, SHAP outperformed T-Explainer; however, our method performed consistently more faithfully to the model’s predictions than the other explainers (columns Top 2 and Top 3).
3H-NN PGI
XAI Top 1 Top 2 Top 3
T-Exp 0.7337 0.7335 0.7334
SHAP 0.6063 0.8402 0.7569
LIME 0.4973 0.4960 0.4956
I-Grad 0.4973 0.4960 0.4956
I×$\times$×Grad 0.4973 0.4960 0.4956
DeepLIFT 0.4973 0.4960 0.4956
Table 3: Faithfulness of XAI methods explaining the 3H-NN predictions on the 4-FT dataset according to the Top k most important features.
Tables 4 and 5 depict the stability of the XRFC and 3H-NN models on the 20-FT dataset. From Table 4, we can see that T-Explainer is superior to TreeSHAP and LIME in the RIS and ROS metrics for both maximum and mean values. TreeSHAP reached a close performance regarding maximum ROS, but T-Explainer is considerably better on the average ROS. Similar results can be observed in Table 5, where T-Explainer outperforms all the other explainers in terms of RIS (3×3$\times$3 × better than Integrated Gradients), with slightly better performance than Integrated Gradients in the mean ROS but outperforming all the other explainers in maximum ROS.
SHAP was the most unstable method, which can be justified in the context of the 20-FT dataset because, in high-dimensional spaces, SHAP uses a random sampling algorithm rather than the deterministic one, leading to unstable explanations [7]. Moreover, SHAP is also prone to suffer from extrapolations [29, 30].
@@
In general, Tables 1, 2, 4, and 5 show that T-Explainer is robust regarding different stability metrics evaluating explanations’ consistency and continuity in controlled synthetic data tasks [42], being among the best performance methods, especially when data dimensionality is high.
Observing the PGI results in Table 6, we have T-Explainer as the most faithful explanation method to the original model predictive behavior for the most important feature (Top 1 column). Enlarging the set of the $k$ most important features (Top 3–9 columns), SHAP presented the best performances; however, T-Explainer consistently outperformed all the other explainers, being the second most faithful explanation method explaining the 3H-NN predictions on the 20-FT synthetic dataset.
3H-NN PGI
XAI Top 1 Top 3 Top 5 Top 7 Top 9
T-Exp 0.5913 0.5738 0.5674 0.5624 0.5669
SHAP 0.5844 0.6540 0.6467 0.6580 0.6724
LIME 0.4586 0.4548 0.4536 0.4536 0.4535
I-Grad 0.4562 0.4540 0.4538 0.4536 0.4536
I×$\times$×Grad 0.5604 0.5621 0.5501 0.5482 0.5504
DeepLIFT 0.5604 0.5621 0.5501 0.5482 0.5504
Table 6: Faithfulness of XAI methods explaining the 3H-NN predictions on the 20-FT dataset according to the Top k most important features.
Table 7 demonstrates XAI explainers’ ability to approximate the original model predictions based on their feature importance attributions (see Equation 5). We used the LAP metric (considering $0.1$ and $0.05$ as tolerance errors, i.e., allowing for $10\%$ and $5\%$ of inaccuracy) to assess the Local Accuracy preservation for feature importance attributions generated to predictions of the 3H-NN model on 4-FT and 20-FT datasets. In the 4-FT columns of Table 7, T-Explainer achieved the best performances for Local Accuracy preservation, either with $0.1$ tolerance error or the more restrictive tolerance of $0.05$. Observe that all the other explainers achieved rates below $0.5$, meaning that except for T-Explainer, all the explanation methods tested violated the Local Accuracy property in more than half of their explanations (remember LAP returns frequency rates of Local Accuracy preservation).
3H-NN LAP ($\epsilon=0.1$) LAP ($\epsilon=0.05$)
XAI 4-FT 20-FT 4-FT 20-FT
T-Exp 0.557 0.494 0.490 0.311
SHAP 0.432 0.281 0.419 0.227
LIME 0.017 0.101 0.010 0.056
I-Grad 0.477 0.166 0.445 0.122
I×$\times$×Grad 0.464 0.377 0.439 0.266
DeepLIFT 0.464 0.377 0.439 0.266
Table 7: Local Accuracy preservation when explaining the 3H-NN predictions on 4-FT and 20-FT datasets.
The 20-FT columns of Table 7 show that the T-Explainer outperformed all the other feature attribution methods regarding Local Accuracy preservation on the 20-FT dataset, considering either the $0.1$ tolerance error or the tighter tolerance of $0.05$. The results in Tables 3, 6, and 7 demonstrate T-Explainer’s capacity for achieving faithful and accurate explanations, while being stable.
In summary, T-Explainer’s performance when dealing with the XRFC classifier, a non-differentiable model, was remarkable for more complex tasks (e.g., explaining predictions for the 20-FT data). Such performances have significant practical implications, suggesting that the T-Explainer might perform well even when applied to generate explanations for non-differentiable machine learning models. The following section presents evaluation results on real-world data.
5.3 Real-world Data
This Section extends the evaluations and comparisons by applying real data from different domains. Specifically, we run experiments using four well-known real-world datasets with distinct properties regarding dimensionality, the presence of categorical attributes, and size.
@@ 5.3 Real-world Data
The Home Equity Line of Credit (HELOC) dataset provided by FICO [25] consists of financial attributes from anonymized applications for home equity lines of credit submitted by $9,871$ real homeowners. The task in the HELOC dataset is to predict whether an applicant has a good or bad risk of repaying the HELOC account within two years. The largest dataset in our study is HIGGS [10], which contains $28$ features about simulated collision events to distinguish between Higgs bosons signals and a background process. The original HIGGS dataset contains $11$ million instances, but we used the $98,050$ instances version available at OpenML [64]. Table 8 summarizes the real datasets.
@@
The experiments with real-world data were conducted using only the neural network models, as some of the explanation models with which T-Explainer is compared are specifically designed for neural networks. Moreover, the neural network models are differentiable, thus meeting the theoretical requirements that support the T-Explainer technique. We further discuss tree-based models in Section 7.
For the experiments with Banknote Authentication and German Credit datasets, we selected the 3H-NN model. From Table 9, one can see that T-Explainer is less stable than Integrated Gradients, but it is the second-best method on average for RIS and ROS (tied with Integrated Gradients for RES).
@@
Analyzing the PGI results in Table 10, we see all the methods tied to a faithful behavior between explanations and model predictions according to the most important feature identified in their explanations. However, T-Explainer’s performance grows when considering the second and third most important features. Only SHAP and T-Explainer achieved a faithfulness higher than $0.5$ for their explanations’ three most important features. In a low-dimensional (four features, see Table 8) set such as the Banknote Authentication, it is possible to conjecture that an explainer with a PGI below $0.5$ is generally missing a highly significant feature in its explanations.
3H-NN PGI
XAI Top 1 Top 2 Top 3
T-Exp 0.4514 0.7304 0.9563
SHAP 0.4514 0.4514 0.4464
LIME 0.4514 0.4514 0.6634
I-Grad 0.4514 0.4514 0.4514
I×$\times$×Grad 0.4514 0.4514 0.4514
DeepLIFT 0.4514 0.4514 0.4514
Table 10: Faithfulness of XAI methods explaining the 3H-NN predictions on the Banknote Authentication dataset according to the Top k most important features.
Table 11 presents the stability tests for the 3H-NN model predictions on the German Credit data. The original version of the dataset has $20$ numerical and categorical attributes. Still, due to its complex categorization system, it is practically impossible to apply it to a machine-learning task directly. We pre-processed the German Credit data to clean it, resulting in a reduced version holding nine features ($4$ numerical and $5$ categorical features, see Table 8). As shown in Table 11, T-Explainer performed as the most stable explainer regarding RIS and maximum ROS perturbations, being the second best for the mean ROS.
@@
Table 12 shows close performances regarding the PGI metric among all feature importance methods for 3H-NN predictions on the German Credit data. LIME, T-Explainer, and SHAP were the most faithful explainers in the incremental verification process of top $k$ features.
3H-NN PGI
XAI Top 1 Top 3 Top 5 Top 7 Top 9
T-Exp 0.8044 0.8164 0.8311 0.8198 0.8178
SHAP 0.8245 0.8193 0.8215 0.8242 0.8240
LIME 0.8396 0.8231 0.8196 0.8184 0.8187
I-Grad 0.8257 0.8182 0.8156 0.8159 0.8159
I×$\times$×Grad 0.8149 0.8212 0.8222 0.8220 0.8220
DeepLIFT 0.8149 0.8212 0.8222 0.8220 0.8220
Table 12: Faithfulness of XAI methods explaining the 3H-NN predictions on the German Credit dataset according to the Top k most important features.
Tables 9 and 11 indicate that gradient-based methods are more stable to input/output perturbations than SHAP and, in some cases, also than LIME. As one can notice, T-Explainer competes quite well with other gradient-based methods regarding stable explanations. While Tables 10 and 12 place T-Explainer between the top feature importance methods for faithful explanations.
HELOC and HIGGS are more robust datasets in terms of dimensionality and size. We selected the 3H-NN and the five-layer neural network models to compare T-Explainer with the other explainers on the real-world datasets. Specifically, we trained the 3H-NN and 5H-128-NN classifiers on HELOC, achieving $71.34\%$ and $73.27\%$ accuracy, respectively. Despite the accuracy values being lower than the ones we achieved before, those performances are in line with the results reported in the literature [16].
Table 13 shows that Integrated Gradients is the most stable explainer for almost all metrics (behind LIME only for mean RIS but with very close numbers). However, we highlight the performance of T-Explainer, especially regarding the metrics’ means. Only Integrated Gradients, LIME, and T-Explainer kept the mean RIS under ten units, which can be taken as virtually the same mean stability to input perturbations. Similar results are observed in mean ROS, where Integrated Gradients, LIME, and T-Explainer achieved the smallest values, being T-Explainer the second best. Table 14 shows the explainers’ stability when applied to the five-layer neural network. T-Explainer was the most stable method, with Integrated Gradients being the second best for mean RIS and LIME as the second best for mean ROS. Only T-Explainer and Integrated Gradients achieved the highest levels of stability in RES.
@@
Tables 15 and 16 present the PGI performances of the explainers. All the methods increased their faithfulness values by including more features to the top $k$ set evaluated, indicating that the HELOC dataset has multiple features significant to the underlying model. For the 3H-NN model, the T-Explainer presented the highest PGI value considering the nine most essential features, i.e., our method could identify nine important features with high fidelity to the model predictions out of more than $20$ predictive features. For the 5H-128-NN model, T-Explainer performed better up to the three most essential features, although our method was the second with higher PGI, after Integrated Gradients, the most faithful explainer to larger top $k$ feature sets.
3H-NN PGI
XAI Top 1 Top 3 Top 5 Top 7 Top 9
T-Exp 0.6360 0.7364 0.8146 0.8595 0.8712
SHAP 0.6309 0.6529 0.7268 0.7740 0.7995
LIME 0.6274 0.6543 0.8350 0.8424 0.8449
I-Grad 0.6379 0.7774 0.8557 0.8614 0.8622
I×$\times$×Grad 0.6379 0.7710 0.8436 0.8539 0.8419
DeepLIFT 0.6379 0.7710 0.8436 0.8539 0.8419
Table 15: Faithfulness of XAI methods explaining the 3H-NN predictions on the HELOC dataset according to the Top k most important features.
5H-128-NN PGI
XAI Top 1 Top 3 Top 5 Top 7 Top 9
T-Exp 0.7028 0.7454 0.7470 0.7689 0.7821
SHAP 0.6373 0.6617 0.6948 0.7329 0.7667
LIME 0.6627 0.7044 0.8022 0.7223 0.7663
I-Grad 0.6763 0.7130 0.7617 0.7923 0.7910
I×$\times$×Grad 0.6506 0.6847 0.7198 0.7445 0.7566
DeepLIFT 0.6506 0.6847 0.7198 0.7445 0.7566
Table 16: Faithfulness of XAI methods explaining the 5H-128-NN predictions on the HELOC dataset according to the Top k most important features.
We trained the 3H-NN and 5H-64-NN classifiers on the HIGGS data, which achieved $69.58\%$ and $66.83\%$ accuracy, respectively, close to values reported in previous works [16]. We included a set of benchmarks using a five-layer neural network here to keep some similarity with the architectures proposed in Baldi et al. [10], which explored the use of deep networks on HIGGS. Tables 17 and 18 show good stability performances of gradient-based feature importance explainers. T-Explainer outperforms them regarding RIS perturbations while being quite competitive regarding ROS and RES metrics. Such results demonstrate T-Explainer’s ability to generate consistent explanations through different levels of model complexity and data dimensionality. Note that SHAP is more unstable than LIME, with SHAP’s claimed advantageous performance predominantly seen in low-dimensional datasets [7].
@@
Tables 19 and 20 present the explainers’ PGI performances on the HIGGS dataset. LIME achieved the highest PGI values when explaining predictions made by the 3H-NN and 5H-64-NN neural networks. However, T-Explainer is the second best regarding faithful explanations when considering larger top $k$ feature sets (see Top 7 and 9 columns).
3H-NN PGI
XAI Top 1 Top 3 Top 5 Top 7 Top 9
T-Exp 0.6322 0.6291 0.6395 0.6297 0.6414
SHAP 0.6355 0.6111 0.6007 0.5920 0.5955
LIME 0.7155 0.7075 0.6841 0.6877 0.6827
I-Grad 0.6380 0.5921 0.5935 0.6071 0.6073
I×$\times$×Grad 0.6704 0.6367 0.6236 0.6231 0.6351
DeepLIFT 0.6704 0.6367 0.6236 0.6231 0.6351
Table 19: Faithfulness of XAI methods explaining the 3H-NN predictions on the HIGGS dataset according to the Top k most important features.
5H-64-NN PGI
XAI Top 1 Top 3 Top 5 Top 7 Top 9
T-Exp 0.6094 0.5923 0.6043 0.6335 0.6357
SHAP 0.5664 0.5838 0.5941 0.6029 0.6292
LIME 0.6768 0.6769 0.6724 0.6545 0.6510
I-Grad 0.6441 0.6074 0.5897 0.5833 0.5916
I×$\times$×Grad 0.6221 0.6109 0.6074 0.6228 0.6215
DeepLIFT 0.6221 0.6109 0.6074 0.6228 0.6215
Table 20: Faithfulness of XAI methods explaining the 5H-64-NN predictions on the HIGGS dataset according to the Top k most important features.
Finally, the experiments above show that T-Explainer generally performs well, clearly outperforming well-known model-agnostic methods such as SHAP (and LIME for most tests). Moreover, T-Explainer also proved quite competitive with model-specific techniques such as gradient-based ones, thus becoming a new and valuable alternative method for explaining black-box models’ predictions.